A commercial real estate developer plans to borrow money to finance an upscale mall in an exclusive area of the city. The developer plans to get a loan that will be repaid with uniform payments of $525,000 beginning in year 2 and ending in year 16. How much will a bank be willing to loan at an interest rate of 13% per year? The bank will be willing to loan the developer a sum of $
Answer #1

Calculate the amount that the bank will be willing to loan, as follows:

P = 525,000(P/A,13%,16-1)/(1+0.13)

= 525,000(P/A,13%,15)/1.13

= 525,000(6.4624)/1.13

= 3,002,488.94(rounded)

=